# CAOS: Conformal Aggregation of One-Shot Predictors

### v1

@@ 1 Introduction
Beyond point predictions, practical deployment in such settings requires reliable uncertainty quantification. Conformal prediction provides a principled framework for constructing prediction sets with finite-sample marginal coverage guarantees (Shafer and Vovk, 2008; Angelopoulos et al., 2024). Applied naively, a common approach is to equip each one-shot predictor with its own conformal prediction set using split conformal calibration. However, data splitting is statistically inefficient in the low-data regimes that motivate one-shot learning (Bashari et al., 2025), and selecting or aggregating predictors adaptively typically violates the assumptions required for conformal validity (Liang et al., 2024).
@@
In this work, we answer this question in the affirmative by introducing Conformal Aggregation of One-Shot Predictors (CAOS), a conformal framework for one-shot prediction. At a high level, CAOS adaptively aggregates nonconformity scores across multiple one-shot predictors while allowing all labeled examples to participate in calibration through a leave-one-out construction. Despite breaking classical exchangeability at the score level, CAOS achieves exact finite-sample marginal coverage guarantees.
Our contributions are threefold: (i) we introduce CAOS, a data-efficient conformal framework for aggregating one-shot predictors without requiring disjoint calibration splits; (ii) we establish exact finite-sample coverage guarantees for CAOS via a novel monotonicity-based reduction to full conformal prediction; and (iii) we demonstrate empirically that CAOS yields substantially smaller prediction sets than split conformal baselines while maintaining reliable coverage in both vision and language one-shot tasks.
@@ 2 Related Work
In-Context Learning (ICL), originally developed in language models (Brown et al., 2020), together with related zero-shot and few-shot prediction paradigms (Bommasani, 2021), has become a standard approach for adapting large pretrained models in both language and vision (Radford et al., 2021; Siméoni et al., 2025). Due to its distribution-free assumptions, conformal prediction (Angelopoulos et al., 2023, 2024) is an attractive framework for uncertainty quantification, as it allows pretrained models to be treated as black boxes while achieving marginal coverage guarantees.
@@
Recent work has applied conformal prediction to large pretrained models in zero-shot, one-shot, and few-shot regimes (Fisch et al., 2021; Park et al., 2023; Fillioux et al., 2024; Quach et al., 2023; Su et al., 2024; Wang et al., 2024; Vishwakarma et al., 2024; Pantelidis et al., 2025; Ye and Wen, 2025; Lin et al., 2025; Tayebati et al., 2025; Silva-Rodríguez et al., 2025a, b; Wang et al., 2025; Xu and Lu, 2025), typically using split conformal calibration for simplicity. In one-shot settings, however, each labeled example induces a distinct predictor, giving rise to a collection of predictors rather than a single fixed model. Existing conformal approaches largely ignore this structure, leaving uncertainty quantification for collections of one-shot predictors underexplored.
@@
While few-shot adaptation is particularly attractive when labeled data are scarce, this regime exacerbates a fundamental trade-off in classical conformal prediction. Split conformal is computationally efficient but statistically inefficient in low-data regimes (Bashari et al., 2025), while full conformal improves statistical efficiency at prohibitive computational cost. Cross-conformal prediction (Vovk, 2015) and related leave-one-out methods (Barber et al., 2021) reuse data more effectively, but typically incur slack in their coverage guarantees due to non-exchangeable scores (Gasparin and Ramdas, 2025), resulting in a tension between efficiency, tractability, and exact validity in the one-shot setting.
When faced with a collection of predictors, a natural response is to either select a single predictor (Liang et al., 2024; Bai and Jin, 2024; Hegazy et al., 2025) or to aggregate predictions across predictors (Gasparin and Ramdas, 2024b; Rivera et al., 2024; Alami et al., 2025). However, most such approaches rely on disjoint data splits or additional calibration data to preserve coverage guarantees, which in turn limits data efficiency in small-sample regimes.
@@
This inefficiency is not incidental, but stems from a central requirement of classical conformal prediction: standard coverage guarantees rely on the exchangeability of conformity scores for the calibration and test examples (Shafer and Vovk, 2008; Angelopoulos et al., 2023, 2024). In settings such as distribution shift (Tibshirani et al., 2019; Barber et al., 2023) or online prediction (Gasparin and Ramdas, 2024a; Sale and Ramdas, 2025), exchangeability is violated at the data level, and valid coverage is typically recovered only at the cost of unavoidable slack terms (Barber et al., 2023). In the one-shot setting considered here, non-exchangeability instead arises algorithmically: aggregation or adaptive selection of predictors based on the observed data breaks exchangeability at the score level, even when the underlying data remain exchangeable, precluding the direct application of classical conformal arguments.
In contrast to the approaches discussed above, CAOS directly addresses the setting that motivates one-shot conformal prediction: small labeled datasets that induce a collection of reference-conditioned predictors whose relevance varies across test inputs. CAOS performs instance-adaptive aggregation of one-shot predictors while allowing all labeled examples to participate in calibration through a leave-one-out construction, thereby avoiding the statistical inefficiencies of data splitting. Although this aggregation breaks exchangeability of the conformity scores, CAOS leverages a monotonicity property  [^1] Interestingly, the theoretical arguments of Jin and Candès (2023) also rely on the monotonicity of the conformity scores, albeit in a different setting (to establish FDR control in multiple selection as opposed to marginal coverage in conformal one-shot prediction) and a different type of monotonicity (monotonicity w.r.t. the labels as opposed to monotonicity w.r.t. a larger reference set). of the aggregated score to recover a sharp $1-\alpha$ finite-sample marginal coverage guarantee via a reduction to full conformal prediction. In this way, CAOS resolves the recurring trade-offs between adaptivity, data efficiency, computational tractability, and exact validity that characterize prior selection- and aggregation-based conformal methods in the one-shot, low-data regime.
@@ 3 Background
where smaller values indicate more conforming predictions. This abstraction allows different one-shot prediction mechanisms to be treated uniformly within the conformal prediction framework. We next describe two concrete instantiations of one-shot prediction and their corresponding nonconformity scores that are used in our experiments.
@@
(see Sec. D.3 for the full prompt template), which conditions the model on a single demonstration.
@@
Large pretrained models may yield strong point predictions when used as one-shot predictors. However, there is no principled way to assess the reliability of these predictions for a new input. In the next section, we introduce split conformal one-shot prediction, which equips each one-shot predictor $\pi_{i}$ with a data-dependent prediction set that enjoys finite-sample marginal coverage guarantees.
@@
Split conformal one-shot prediction yields a valid prediction set for each reference example, but the resulting sets vary widely in size and usefulness, reflecting heterogeneity in one-shot predictor quality. Ideally, predictions would adaptively select or combine the most relevant reference examples for each test input. This creates a basic tension in the one-shot setting: predictive performance benefits from reusing all labeled examples, while conformal calibration restricts such reuse. We resolve this tension with Conformal Aggregation of One-Shot Predictors (CAOS), a conformal framework that aggregates one-shot predictors while allowing all labeled examples to participate in calibration.
@@ 4 Method
Problem setting and goal.  We consider a one-shot prediction setting with $n$ labeled examples $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Each labeled example induces a one-shot predictor $\pi_{i}$ with an associated nonconformity score $s_{\pi_{i}}$. Our goal is to construct prediction sets $\hat{C}_{\mathrm{caos}}(X_{n+1})$ that satisfy finite-sample marginal coverage guarantees while remaining as small as possible. Smaller prediction sets correspond to more informative predictions and reduced predictive uncertainty.
No single reference example is uniformly optimal: the informativeness of one-shot predictors varies substantially across reference examples and may depend on the test input. The central challenge is therefore to design a method that (i) aggregates information across multiple one-shot predictors, (ii) calibrates the resulting aggregate using the same limited labeled data, and (iii) retains finite-sample coverage guarantees. We address this challenge with CAOS, a leave-one-out conformal aggregation strategy that leverages all labeled data for both aggregation and calibration.
@@
$s_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n}\bigr)\;:=\;\frac{1}{k}\,\Sigma^{k}_{\min}(\mathcal{A}_{\mathcal{D}_{n}}\!\left(X_{n+1},y\right)),$ (13)
where $k$ is a small constant (we use $k=3$ in all experiments). This score is small when several reference examples strongly support the candidate output $y$, and large otherwise. We explicitly indicate the dependence on $\mathcal{D}_{n}$, since unlike split conformal prediction, CAOS does not maintain disjoint reference and calibration sets.
@@
The CAOS nonconformity score is
$S_{\mathrm{caos}}^{i}\;:=\;\frac{1}{k}\,\Sigma^{k}_{\min}(\mathcal{A}_{\mathcal{D}_{n}^{-i}}\!\left(X_{i},Y_{i}\right)).$ (14)
Collecting these scores over all $i=1,\ldots,n$, we compute the empirical quantile
@@
CAOS’ coverage guarantees.  The leave-one-out calibration procedure treats test points differently from calibration data and therefore breaks the classical exchangeability assumptions used in conformal prediction. To establish valid coverage guarantees, we make the following assumption.
Assumption 4.1 (Self-score optimality).
For any labeled example $(X_{i},Y_{i})$ in any dataset $\mathcal{D}_{n}$, the one-shot nonconformity score induced by using $(X_{i},Y_{i})$ as its own reference is minimal among all one-shot scores for that target, i.e.,
$s_{\pi_{i}}(X_{i},Y_{i})\;\leq\;s_{\pi_{j}}(X_{i},Y_{i})\quad\text{for all }(X_{j},Y_{j})\in\mathcal{D}_{n}.$ (17)
This assumption reflects the natural property that a labeled example is maximally informative about itself, which holds in common one-shot settings such as patch similarity.
@@
Assume that the labeled data $\mathcal{D}_{n}$ and the test example $(X_{n+1},Y_{n+1})$ are exchangeable and that Assumption 4.1 holds. Then the CAOS prediction set defined in Eqn. 16 satisfies the finite-sample marginal coverage guarantee
@@
The proof is nontrivial because the CAOS nonconformity scores are not exchangeable. In Sec. 5, we establish Thrm. 4.2 via a reduction to a theoretical full conformal variant of CAOS.
Algorithm 1 CAOS Prediction
[…]
@@ 5 Theory: Coverage Guarantees for CAOS
This Section establishes Thrm. 4.2. The proof proceeds via a reduction to a theoretical full conformal version of CAOS. The argument relies on two key observations: (i) a monotonicity property of the CAOS nonconformity score with respect to the reference dataset, and (ii) a set inclusion result showing that CAOS prediction sets contain those produced by full conformal prediction. Together, these imply that CAOS inherits the finite-sample marginal coverage guarantee of full conformal prediction, despite using a computationally cheaper scoring rule that does not yield exchangeable scores.
@@ 5.1 Theoretical Construction of Full CAOS
Full conformal prediction is a rigorous but impractical version of conformal prediction. Split conformal prediction as described in Sec. 3, is often preferred due to its computational efficiency, which requires the scores to be calibrated only once, whereas full conformal prediction requires calibrating the threshold for each test example for each possible prediction label separately. [^2] In classical machine learning settings the underlying predictive model needs to also be retrained as many times, but in the one-shot setting considered here, there is at least no training involved. However, the computational efficiency of split conformal comes at the cost of statistical efficiency: while full conformal (and CAOS) use all of the labeled examples for calibration, split conformal can only use a fraction, e.g. $n/2$.
In this section, we describe a full conformal version of CAOS. This algorithm will never be run in practice due to the large computational cost. It is only needed for the theory to establish a rigorous coverage guarantee for CAOS as presented in Sec. 4.
@@
$\tilde{S}_{\mathrm{full}}^{i}$ $=\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n+1}^{y}\bigr)$ (19)
$=\frac{1}{k}\bigl(\Sigma^{k+1}_{\min}(\mathcal{A}_{\mathcal{D}_{n+1}^{y}}\!\left(X_{i},Y_{i}\right))-s_{\pi_{i}}(X_{i},Y_{i})\bigr).$
This definition differs slightly from the leave-one-out score $s_{\mathrm{caos}}$ defined in Eqn. 14. However, it is deliberately chosen to i) establish exchangeability of full CAOS and ii) such that, under Assump. 4.1, the test-time nonconformity scores of CAOS and full CAOS coincide, i.e.,
$\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n+1}^{y}\bigr)=s_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n}\bigr).$ (20)
This equivalence (proof in Sec. C.1) will play a key role in the set-inclusion argument used to establish coverage for CAOS. Based on the calibration scores $\tilde{S}_{\mathrm{full}}^{i}$, full CAOS requires a separate threshold for each test input $y$
@@
To establish conformal validity, full conformal prediction requires the scores to by symmetric.
@@
Proof.
The operator $\Sigma^{k+1}_{\min}$ depends only on the multiset of its inputs and is therefore symmetric. Moreover, the subtracted term in Eqn. 19 corresponds to the one-shot score induced by the target example itself and is unaffected by permutations of the remaining elements of $\mathcal{D}$. It follows that $\tilde{s}_{\mathrm{caos}}((x,y);\mathcal{D})=\tilde{s}_{\mathrm{caos}}((x,y);\mathcal{D}_{\sigma}).$
By Lemma 5.1 and standard full conformal arguments (Angelopoulos et al., 2024), the prediction set defined in Eqn. 22 achieves finite-sample marginal coverage guarantees at the target level. We prove next that these guarantees extend to CAOS.
@@ 5.2 Conformal Validity of CAOS
We now prove Thrm. 4.2 by comparing the two constructions of CAOS: the leave-one-out calibrated prediction set defined in Eqn. 16 and the theoretical full conformal construction defined in Eqn. 22. The proof leverages the conformal validity of full CAOS via two steps: (i) a monotonicity lemma showing that the CAOS nonconformity scores can only improve as the reference set grows, and (ii) a comparison lemma relating leave-one-out CAOS scores to their full conformal analogues. Together, these yield the set inclusion $\hat{C}_{\mathrm{full}}(X_{n+1})\subseteq\hat{C}_{\mathrm{caos}}(X_{n+1})$, which implies coverage.
@@
$\Sigma^{k}_{\min}(\mathcal{A}^{\prime})=\sum_{j=1}^{k}a^{\prime}_{(j)}\;\leq\;\sum_{j=1}^{k}a_{(j)}=\Sigma^{k}_{\min}(\mathcal{A}).$
Dividing by $k$ yields $s_{\mathrm{caos}}((x,y);\mathcal{D}^{\prime})\leq s_{\mathrm{caos}}((x,y);\mathcal{D}),$ as claimed. ∎
We can now use this Lemma to compare the calibration scores of CAOS and full CAOS.
Lemma 5.3 (Comparison of CAOS and full CAOS scores).
Fix a candidate label $y\in\mathcal{Y}$ and the augmented dataset $\mathcal{D}_{n+1}^{y}=\mathcal{D}_{n}\cup\{(X_{n+1},y)\}$. Then for all $i=1,\ldots,n$,
Proof.
By the same argument in App. C.1 that uses Assum. 4.1 (self-score optimality) to establish the equivalence of test time scores (Eqn. 20), we have
$\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)$
$=s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,$ $\mathcal{D}_{n}^{-i}\cup\{X_{n+1},y\}\bigr).$ (24)
Since $\mathcal{D}_{n}^{-i}\cup\{X_{n+1},y\}$ contains $\mathcal{D}_{n}^{-i}$, the claimed inequality follows from the monotonicity lemma (Lemma 5.2). ∎
@@
Let $\hat{C}_{\mathrm{full}}(X_{n+1})$ be the full conformal CAOS set defined in Eqn. 22, and let $\hat{C}_{\mathrm{caos}}(X_{n+1})$ be the CAOS set defined in Eqn. 16. Then for any test input $X_{n+1}$,
@@
Let $y\in\hat{C}_{\mathrm{full}}(X_{n+1})$. By definition of the prediction set (Eqn. 22), $y\leq\hat{q}_{\mathrm{full}}^{y}$. By Lemma 5.3, $\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)\leq s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr)$ for all calibration indices $i=1,\cdots,n$. These calibration scores are used to establish $\hat{q}_{\mathrm{caos}}$ and $\hat{q}_{\mathrm{full}}^{y}$ via Eqns. 15 and 21 respectively. Since the quantile function $\operatorname{Quantile}\!\left(\cdot\,;\,(1-\alpha)(1+1/n)\right)$ is monotonic in its first argument, the resulting thresholds obey $\hat{q}_{\mathrm{full}}^{y}\leq\hat{q}_{\mathrm{caos}}$, which means that $y\leq\hat{q}_{\mathrm{caos}}$. Therefore, by Eqn. 16, $y\in\hat{C}_{\mathrm{caos}}(X_{n+1})$. ∎
Proof of Thrm. 4.2.
By the symmetry of $\tilde{s}_{\mathrm{caos}}\bigl((x,y)\,;\,\mathcal{D}_{n+1}^{y}\bigr)$ (Lemma 5.1) and the assumed exchangeability of samples from the augmented dataset $\mathcal{D}_{n+1}^{y}$, standard full conformal arguments (Angelopoulos et al., 2024) imply finite-sample marginal coverage:
$\Pr\!\left(Y_{n+1}\in\hat{\mathcal{C}}_{\mathrm{full}}(X_{n+1})\right)\;\geq\;1-\alpha.$ (26)
Next, by the set inclusion argument (Corr. 5.4), the event $\{Y_{n+1}\in\hat{\mathcal{C}}_{\mathrm{full}}(X_{n+1})\}$ implies the event $\{Y_{n+1}\in\hat{C}_{\mathrm{caos}}(X_{n+1})\}$, and hence
$\Pr\!\left(Y_{n+1}\in\hat{C}_{\mathrm{caos}}(X_{n+1})\right)\;$ $\geq\;\Pr\!\left(Y_{n+1}\in\hat{\mathcal{C}}_{\mathrm{full}}(X_{n+1})\right)\;$
Together with Eqn. 26, this establishes Thrm. 4.2. ∎
Remarks.  The CAOS coverage guarantee applies to nonconformity scores that satisfy two structural properties: (i) symmetry with respect to permutations of the reference data, and (ii) monotonicity with respect to the inclusion of additional examples. Score functions that violate these conditions, such as scores involving parameters retrained on the augmented dataset, fall outside the scope of our theory.

### v2

@@ 1 Introduction
Beyond point predictions, practical deployment in such settings requires reliable uncertainty quantification. Conformal prediction provides a principled framework for constructing prediction sets with finite-sample marginal coverage guarantees (Shafer & Vovk, 2008; Angelopoulos et al., 2024). Applied naively, a common approach is to equip each one-shot predictor with its own conformal prediction set using split conformal calibration. However, data splitting is statistically inefficient in the low-data regimes that motivate one-shot learning (Bashari et al., 2025), and selecting or aggregating predictors adaptively typically violates the assumptions required for conformal validity (Liang et al., 2024).
@@
In this work, we answer this question in the affirmative by introducing Conformal Aggregation of One-Shot Predictors (CAOS), a conformal framework for one-shot prediction. At a high level, CAOS adaptively aggregates nonconformity scores across multiple one-shot predictors while allowing all labeled examples to participate in calibration through a leave-one-out construction. Despite breaking classical exchangeability at the score level, CAOS achieves valid finite-sample marginal coverage guarantees.
Our contributions are threefold: (i) we introduce CAOS, a data-efficient conformal framework for aggregating one-shot predictors without requiring disjoint calibration splits; (ii) we establish finite-sample marginal coverage guarantees for CAOS via a novel monotonicity-based reduction to full conformal prediction; and (iii) we demonstrate empirically that CAOS yields substantially smaller prediction sets than split conformal baselines while maintaining reliable coverage in both vision and language one-shot tasks.
@@ 2 Related Work
In-Context Learning (ICL), first popularized for language models (Brown et al., 2020) and closely related to zero-shot and few-shot prediction paradigms (Bommasani, 2021), is now a standard mechanism for adapting large pretrained models in both language and vision (Radford et al., 2021; Siméoni et al., 2025). In parallel, conformal prediction provides a distribution-free framework for uncertainty quantification with finite-sample marginal coverage guarantees (Shafer & Vovk, 2008; Angelopoulos et al., 2023, 2024), and is particularly appealing in foundation-model settings because it can treat the underlying predictor as a black box.
@@
A growing body of work applies conformal prediction to foundation models in zero-shot, one-shot, and few-shot regimes (Fisch et al., 2021; Quach et al., 2023; Su et al., 2024; Wang et al., 2024; Fillioux et al., 2024); see also (Park et al., 2023; Vishwakarma et al., 2024; Pantelidis et al., 2025; Ye & Wen, 2025; Lin et al., 2025; Tayebati et al., 2025; Silva-Rodríguez et al., 2025a, b; Wang et al., 2025; Xu & Lu, 2025). Many of these methods adopt split conformal calibration for simplicity and modularity. In the one-shot setting, each labeled example induces its own predictor, yielding a collection of predictors rather than a single model. Existing conformal approaches typically do not address uncertainty quantification for collections of one-shot predictors whose relevance may vary across test inputs.
@@
The one-shot, low-data regime also sharpens a classical trade-off in conformal prediction between statistical and computational efficiency. Split conformal is computationally efficient but can be statistically inefficient when labeled data are scarce (Bashari et al., 2025), while full conformal can be more statistically efficient but is often computationally prohibitive. Data-reuse schemes such as cross-conformal prediction (Vovk, 2015) and related leave-one-out approaches (Barber et al., 2021) improve data efficiency, but their validity analyses frequently require additional arguments and may introduce slack when the resulting conformity scores are not exchangeable (Gasparin & Ramdas, 2025). These considerations motivate methods that reuse labeled data while retaining sharp finite-sample guarantees.
When confronted with a pool of predictors, a natural strategy is to either select a single predictor (Liang et al., 2024; Bai & Jin, 2024; Hegazy et al., 2025) or aggregate predictors or their uncertainty sets (Gasparin & Ramdas, 2024b; Rivera et al., 2024; Alami et al., 2025). However, many selection- and aggregation-based conformal approaches preserve validity by relying on disjoint data splits, extra calibration data, or conservative corrections.
@@
These limitations trace to a central requirement in classical conformal prediction: standard finite-sample guarantees rely on exchangeability of the calibration and test conformity scores (Shafer & Vovk, 2008; Angelopoulos et al., 2023, 2024). When exchangeability is violated at the data level, for example under distribution shift (Tibshirani et al., 2019; Barber et al., 2023) or in online prediction (Gasparin & Ramdas, 2024a; Sale & Ramdas, 2025), validity is often recovered via additional assumptions or correction factors that can loosen guarantees (Barber et al., 2023). In the one-shot setting studied here, non-exchangeability instead arises algorithmically, because instance-adaptive selection or aggregation over a pool of reference-conditioned predictors breaks exchangeability at the score level, preventing a direct application of classical conformal arguments. We address these challenges with CAOS, which performs instance-adaptive aggregation with leave-one-out calibration while retaining $1-\alpha$ finite-sample marginal coverage.
@@ 3 Background
where smaller values indicate more conforming predictions. These scores will allow us to convert one-shot predictions into conformal prediction sets. We next describe two instantiations of one-shot prediction and their corresponding nonconformity scores that are used in our experiments.
@@
(see Sec. C.3 for the full prompt template), which conditions the model on a single demonstration.
@@
Large pretrained models may yield strong point predictions when used as one-shot predictors. However, there is also a need to assess the reliability of these predictions for a new input. In the next section, we introduce split conformal one-shot prediction, which equips each one-shot predictor $\pi_{i}$ with a data-dependent prediction set that enjoys finite-sample marginal coverage guarantees.
@@
Split conformal one-shot prediction yields a valid prediction set for each reference example, but the resulting sets vary widely in size and usefulness, reflecting heterogeneity in one-shot predictor quality. Ideally, predictions would adaptively select or combine the most relevant reference examples for each test input. This creates a basic tension in the one-shot setting: predictive performance benefits from using all labeled examples, while conformal calibration restricts such reuse. We resolve this tension with Conformal Aggregation of One-Shot Predictors (CAOS), a conformal framework that aggregates one-shot predictors while allowing all labeled examples to participate in calibration.
@@ 4 Method
Problem setting and goal.  We consider a one-shot prediction setting with $n$ labeled examples $\mathcal{D}_{n}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Each labeled example induces a one-shot predictor $\pi_{i}$ with an associated nonconformity score $s_{\pi_{i}}$. Our goal is to construct prediction sets $\hat{C}_{\mathrm{caos}}(X_{n+1})$ that are small but satisfy finite-sample marginal coverage guarantees. Smaller prediction sets correspond to more informative predictions and reduced predictive uncertainty.
No single reference example is uniformly optimal: the informativeness of one-shot predictors varies substantially across reference examples and may depend on the test input. The central challenge is therefore to design a method that (i) aggregates information across multiple one-shot predictors, (ii) calibrates the resulting aggregate using the same limited labeled data, and (iii) retains finite-sample coverage guarantees. We address this challenge with CAOS, a simple aggregation strategy that thanks to a novel theoretical argument allows all labeled data to participate in aggregation and calibration while maintaining valid coverage.
Algorithm 1 CAOS Prediction
[…]
@@
$s_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n}\bigr)\;:=\;\Sigma^{k}_{\min}(\mathcal{A}_{\mathcal{D}_{n}}\!\left(X_{n+1},y\right)),$ (13)
where $k<n$ is a small constant (we fix $k=3$; App. D studies sensitivity to $k$). This score is small when several reference examples strongly support the candidate output $y$, and large otherwise. We explicitly indicate the dependence on $\mathcal{D}_{n}$, since unlike split conformal prediction, CAOS does not maintain disjoint reference and calibration sets.
@@
The CAOS calibration scores are
$s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr)\;=\;\Sigma^{k}_{\min}(\mathcal{A}_{\mathcal{D}_{n}^{-i}}\!\left(X_{i},Y_{i}\right)).$ (14)
Collecting these scores $S_{\mathrm{caos}}^{i}:=s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr)$ over all $i=1,\ldots,n$, we compute the empirical quantile
@@
CAOS’ coverage guarantees.  The leave-one-out calibration procedure treats test points differently from calibration data and therefore breaks the classical exchangeability assumptions used in conformal prediction. Yet, we will establish the following result for CAOS.
@@
Assume that $\mathcal{D}_{n}$ and the test example $(X_{n+1},Y_{n+1})$ are exchangeable. Then the CAOS prediction set defined in Eqn. 16 satisfies the finite-sample marginal coverage guarantee
@@
The proof is nontrivial because the CAOS scores are not exchangeable. In Sec. 5, we establish Thrm. 4.1 via a reduction to a theoretical full conformal variant of CAOS.
@@ 5 Theory: Coverage Guarantees for CAOS
This Section establishes Thrm. 4.1. The proof proceeds via a reduction to a theoretical full conformal version of CAOS. The argument relies on two key observations: (i) a monotonicity property of the CAOS nonconformity score with respect to the reference dataset, and (ii) a set inclusion result showing that CAOS prediction sets contain those produced by full conformal prediction. Together, these imply that CAOS inherits the finite-sample marginal coverage guarantee of full conformal prediction, despite using a computationally cheaper scoring rule that does not yield exchangeable scores.
@@ 5.1 Theoretical Construction of Full CAOS
Full conformal prediction provides finite-sample marginal coverage but is computationally prohibitive. It requires recalibration for each possible label for each test input. We present a full-conformal version of CAOS not intended for practical use, solely as an analytical tool to establish CAOS’ coverage guarantees.
@@
$\tilde{S}_{\mathrm{full}}^{i}$ $=\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)$ (18)
$=\Sigma^{k}_{\min}\bigl(\mathcal{A}_{\mathcal{D}_{n+1}^{y}}\!\left(X_{i},Y_{i}\right)-\{s_{\pi_{i}}(X_{i},Y_{i})\}\bigr),$
where - denotes the multiset difference as defined in Eqn. 30, App. F. The full CAOS scores differ slightly from the leave-one-out definition of $s_{\mathrm{caos}}$ in Eqn. 14. However, it is deliberately chosen to i) establish exchangeability of full CAOS and ii) such that, the relationship between the scores of CAOS and full CAOS implies that the full conformal coverage guarantees also extend to CAOS.
Based on the calibration scores $\tilde{S}_{\mathrm{full}}^{i}$, full CAOS requires a separate threshold for each test input $y$
@@
To establish conformal validity, full conformal prediction requires the scores to be symmetric.
@@
(Proof in Appendix, Sec. B.1.) By Lemma 5.1 and standard full conformal arguments (Angelopoulos et al., 2024), the prediction set defined in Eqn. 20 achieves finite-sample marginal coverage guarantees at the target level. We prove next that these guarantees extend to CAOS.
@@ 5.2 Conformal Validity of CAOS
We now prove Thrm. 4.1 by comparing the two constructions of CAOS: the leave-one-out calibrated prediction set defined in Eqn. 16 and the theoretical full conformal construction defined in Eqn. 20. The proof leverages the conformal validity of full CAOS via two steps: (i) a monotonicity lemma showing that the CAOS nonconformity scores can only improve as the reference set grows, and (ii) a comparison lemma relating leave-one-out CAOS scores to their full conformal analogues. Together, these yield the set inclusion $\hat{C}_{\mathrm{full}}(X_{n+1})\subseteq\hat{C}_{\mathrm{caos}}(X_{n+1})$, which implies coverage.
The following monotonicity lemma formalizes the intuition, that adding more reference examples will only affect the $\Sigma^{k}_{\min}$ aggregation if it introduces small scores, and in this case, it can only improve the aggregated score.
@@
Lemma 5.2 (proof in App. B.2) enables a comparison of the CAOS calibration scores with full CAOS.
Lemma 5.3 (Comparison of CAOS and full CAOS scores).
Given a candidate label $y\in\mathcal{Y}$ and the augmented dataset $\mathcal{D}_{n+1}^{y}=\mathcal{D}_{n}\cup\{(X_{n+1},y)\}$, the test scores are equivalent
(Test score equivalence)
while for the calibration scores indexed by $i=1,\ldots,n$,
$\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)\;\leq\;s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr),$ (22)
$\text{(Calibration score dominance)}.$
@@
Both statements follow directly from the definitions of the CAOS and full CAOS scores. For the test point $(X_{n+1},y)$, the full CAOS construction removes the self-score, and the remaining reference set coincides with $D_{n}$, yielding the test score equivalence. For calibration points, removing the self-score $i$ from the augmented dataset leaves a reference set that strictly contains $D_{n}^{-i}$, so the inequality in Eqn. 22 follows from the monotonicity of the $\Sigma_{k}^{\min}$ aggregation (Lemma 5.2). Complete derivations for both cases are provided in Appendix B.3. ∎
@@
Let $\hat{C}_{\mathrm{full}}(X_{n+1})$ be the full conformal CAOS set defined in Eqn. 20, and let $\hat{C}_{\mathrm{caos}}(X_{n+1})$ be the CAOS set defined in Eqn. 16. Then for any test input $X_{n+1}$,
@@
Let $y\in\hat{C}_{\mathrm{full}}(X_{n+1})$. Using Lemma 5.3, denote $s:=\tilde{s}_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n+1}^{y}\bigr)=s_{\mathrm{caos}}\bigl((X_{n+1},y)\,;\,\mathcal{D}_{n}\bigr)$. By definition of the prediction set (Eqn. 20), $s\leq\hat{q}_{\mathrm{full}}^{y}$. By Lemma 5.3, $\tilde{s}_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n+1}^{y}\bigr)\leq s_{\mathrm{caos}}\bigl((X_{i},Y_{i})\,;\,\mathcal{D}_{n}^{-i}\bigr)$ for all calibration indices. These calibration scores are used to establish $\hat{q}_{\mathrm{caos}}$ and $\hat{q}_{\mathrm{full}}^{y}$ via Eqns. 15 and 19. Since $\operatorname{Quantile}\!\left(\cdot\,;\,(1-\alpha)(1+1/n)\right)$ is monotonic in its first argument, the resulting thresholds obey $\hat{q}_{\mathrm{full}}^{y}\leq\hat{q}_{\mathrm{caos}}$, which means that $s\leq\hat{q}_{\mathrm{caos}}$. Therefore, by Eqn. 16, $y\in\hat{C}_{\mathrm{caos}}(X_{n+1})$. ∎
Proof of Thrm. 4.1.
We established $\hat{\mathcal{C}}_{\mathrm{full}}(X_{n+1})\subseteq\hat{C}_{\mathrm{caos}}(X_{n+1})$ for any $X_{n+1}$ via Cor. 5.4, so it suffices to show that $\hat{\mathcal{C}}_{\mathrm{full}}$ has $1-\alpha$ marginal coverage. This follows from standard full conformal arguments, using the symmetry of $\tilde{s}_{\mathrm{caos}}\bigl((x,y)\,;\,\mathcal{D}_{n+1}^{y}\bigr)$ (Lemma 5.1) and the assumed exchangeability of samples from $\mathcal{D}_{n+1}^{y}$. ∎
Remarks.  The CAOS coverage guarantee applies to nonconformity scores that satisfy two structural properties: (i) symmetry with respect to permutations of the reference data, and (ii) monotonicity with respect to the inclusion of additional examples. Score functions that violate these conditions, such as scores involving fine-tuning or hyperparameter selection, fall outside the scope of our theory.